# Infinite-Width Limit of a Single Attention Layer: Analysis via Tensor Programs

### v2

@@ 1 Introduction
A useful approach to understanding the complex probabilistic behavior of neural networks is through the study of parameter distributions in the infinite-width limit. Notable examples include the neural network Gaussian process (NNGP) LBN+ (18); HBSDN (20), which approximates the limit of stochastic parameter distributions with Gaussian processes; the neural tangent kernel (NTK) JGH (18), which represents the model near the initial value with kernel functions; mean field theory, which describes the update of parameter distributions MMN (18); and the Tensor Program, developed by Yan19a ; Yan20a ; YL (21); Yan20b ; YH (21); YL (23); YHB+ (21); YYZH (24), which is a general probabilistic analytical framework that unifies the representation of the infinite-width limit for a wide range of neural architectures and multiple layers. These methods provide probabilistic models that closely approximate the complex phenomena of neural networks.
One challenge in the studies is to properly represent the attention mechanisms used in Transformers VSP+ (17), which frequently appear in recent large-scale architectures DCLT (19); AAA+ (23); ZZL+ (23). Unlike ordinary multi-layer perceptrons, an attention layer has interactions between query and key variables by multiplication, which makes the infinite-width limit distribution considerably more complex as shown in HBSDN (20), for example. To avoid this difficulty, the existing studies have mainly limited themselves to two special settings to compute the parameter distribution of the limit: (i) Infinite-head regime: HBSDN (20) considers the NNGP for the attention mechanism with infinite heads, resulting in a Gaussian approximation. (ii) $1/n$-scaling regime: the Tensor Programs Yan19b approximate the multiplication by changing the scale of the input variables for attention layers from $1/\sqrt{n}$ to $1/n$, where $n$ is a dimension of variables. However, these simplifications compromise the expressiveness and structure of actual attention mechanisms (and consequently, of Transformers). Specifically, the limit distribution at the infinite-head is not an effective approximation of the actual attention mechanisms because it differs significantly from the finite-head case. Also, the $1/n$-scaling regime makes all similarity scores converge to zero in the infinite-width limit, making the model equivalent to not measuring similarity between key and query vectors. Therefore, the limit parameter distribution of the attention mechanism is still under development.
@@ 1.1 Related Works
The concentration of measures in neural networks plays a fundamental role in both theoretical analysis and practical applications of machine learning. Initially, studies in this area aimed to characterize the feedforward signal propagation in wide neural networks with random weights, inspired by statistical mechanics Ama (77); SCS (88). The outputs of such random neural networks converge to a Gaussian process, and the computation of signal propagation reduces to the composition of kernel functions of Gaussian processes utilized in machine learning Nea (96); Wil (96); DFS (16); LBN+ (18); dGMHR+ (18). They are often referred to as the NNGP. Since the NNGP kernel captures intrinsic inductive biases of architectures, its prediction performance correlates well with trained networks across different architectures LBN+ (18). Moreover, one can interpret NNGP as a network with random initialization of optimization, which naturally leads to quantitative insight into desirable weight scales to avoid exploding/vanishing signal and gradient problems PLR+ (16); SGGSD (17).
@@
Depending on the network architecture, various types of kernels can be obtained, not only for fully-connected neural networks, but also convolutional neural networks (CNNs) NXL+ (19), skip connections YS (17), naive or gated recurrent neural networks (RNNs) CPS (18); Yan19b , and more YPR+ (19); GHLG (23). While some classical works assumed random Gaussian weights generated in an i.i.d. manner, recent research has shown that the same NNGP can be derived even in networks with non-Gaussian GY (22) or weakly correlated weights SNT (24). The NNGP can also be obtained for a network with a narrow bottleneck layer sandwiched between wide layers APH (20).
@@
Two pioneering studies have investigated the NNGP of self-attention layers Yan19b ; HBSDN (20). Initially, Yan19b pointed out that an unconventional scaling factor of $1/n$ in the softmax function enables straightforward NNGP evaluation. Greg Yang has introduced the theoretical framework of Tensor Programs Yan19a ; Yan20a ; Yan20b , systematically composing kernel functions, including NNGPs, for modern neural networks. Transformers with the $1/n$-scaling fall within the scope of the applicability of Tensor Programs. However, the $1/n$-scaling is rarely used in theoretical or practical contexts, leaving the more realistic $1/\sqrt{n}$-scaling unresolved. To attack this problem, HBSDN (20) analyzed self-attention by varying the number of heads. They numerically demonstrated non-Gaussian behavior emerging in the single-head case due to stochasticity and correlations within the attention matrix, even at infinite embedding dimensions. They further showed that if we take the infinite limit of the number of heads, Gaussian behavior emerges in the self-attention output and defines an NNGP kernel termed infinite attention. Although this infinite attention empirically improved performance on certain NNGP regression benchmarks, its suitability as a theoretical foundation for realistic self-attention remains uncertain. This is because practical attention implementations typically use only 1-128 heads EXW+ (24), far fewer than the embedding dimension.
@@
Beyond the classical NNGP and Tensor Programs analyses, several recent works have examined Transformers under the standard $1/\sqrt{n}$ scaling and related asymptotic or dynamical regimes. DYZ (23) analyzed Transformers by tracking the first two moments (the kernel) of the signal to characterize propagation at initialization and during training. CNQG (24) applied mean-field theory to characterize the edge of chaos via forward and backward signal propagation, assuming Gaussianity of the QK product. BCP (24) employed dynamical mean field theory to study training dynamics under various infinite limits, including infinite width, heads, and depth, identifying parameterizations that ensure stable feature learning over time. NLL+ (23) modeled signal evolution with a stochastic differential equation, which requires modifying the softmax function.
@@ 2.1 Notation and Setup of Neural Networks
We define notation for standard neural networks, excluding the attention mechanism. In particular, we adopt notations inspired by the framework of Tensor Programs Yan19a ; Yan19b , which allow us to describe a broad class of neural network architectures.
@@
This type of Tensor Program is called Netsor, and it covers a broad range of neural network architectures, including a perceptron layer, a convolutional layer, a recurrent layer, and many others Yan19a ; Yan19b .
@@
Here, the common scaling $1/\sqrt{n}$ in the definition of $p_{i,j}^{(a)}$ is a key issue in this study. [^1] Our analysis focuses on the $1/\sqrt{n}$-scaling, a choice motivated by its prevalence in both practical Transformer implementations and the theoretical literature (e.g., HBSDN (20); DYZ (23); CNQG (24)). While some analyses have raised questions about its stability YHB+ (21); BCP (24), the precise conditions for stable training remain an active area of research and may depend on factors often abstracted away in simplified limits, such as the number of tokens or data-specific statistics. For instance, recent work has shown that data statistics can significantly alter optimal scaling rules HL (25). Applying the row-wise SoftMax to the score matrix, we define
@@
This setup of weight matrices is common in practice (see GB (10) for summary) and also identical to that in the existing Tensor Programs Yan19a ; Yan19b ; Yan20b .
@@ 2.2 Distribution of Attention Mechanism in Previous Setup
We discuss the existing challenges on characterizing the output distribution of the attention layer, which exhibits non-Gaussian behavior. To derive this distribution, two regimes are typically made: the choice of scaling in the similarity computation, and the number of heads.
@@
Yan19b studies an attention layer whose scaling term $1/\sqrt{n}$ in the scaled dot-product of Eq. (2) is replaced by $1/n$. In this regime, the Tensor Programs framework can show that the outputs of the attention layer converge to a Gaussian distribution: as $n\to\infty$, it holds that
@@
with some $\kappa>0$ (see Appendix A and Theorem E.8 in Yan19b for details). Intuitively, in the $1/n$-scaling regime, the dot-product score $p_{i,j}^{(a)}$ converges to 0 for every pair $(i,j)\in[s]^{2}$ and head $a\in[H]$, which simplifies the non-Gaussian behavior of the attention outputs.
@@
HBSDN (20) considers the regime in which many dimensionality diverge to infinite, then shows the convergence of the attention outputs to Gaussian. Specifically, it is shown that as $n,H\to\infty$, it holds that
@@
with some $\kappa^{\prime}>0$ (see Theorem 3 in HBSDN (20)). The variance $(\kappa^{\prime})^{2}$ is described by a covariance of the nonlinearly transformed dot-product score $p_{i,j}^{(a)}$. In the result, by letting the number of heads $H$ grow to infinite, the complex non-Gaussian effects from the dot-product score are smoothed out, yielding a convenient Gaussian distribution.
While these approximations are analytically appealing, they have notable limitations. In practice, Transformers typically use $1/\sqrt{n}$-scaling (see Equation (1) in VSP+ (17)) and a finite number of heads, resulting in the frequently observed non-Gaussian behaviors of attention layers. Although the above regimes simplify the behavior into a Gaussian form for tractability, a more precise theory is needed to capture the true behavior of attention layers in practice.
@@ 3.1 Limiting Distribution
In Definition 3.1, the formulations for the limiting random variables $Z^{h}$ associated with initial vectors, MatMul outputs, and Nonlin outputs are developed by the existing Tensor Programs framework (Yan19b, ). Specifically, the limiting distribution of initial vectors follows that given in Section 2.1; variables generated by MatMul converge to Gaussian distribution; and variables generated by Nonlin are nonlinear transformations of the corresponding limiting distributions.
The distinct aspect of our analysis lies in the treatment of the limiting distribution for the scalar dot-products $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, which represent pre-softmax attention scores. While the marginal Gaussian distribution of these scores is consistent with that in the infinite-head limit (see Theorem 3 in HBSDN (20)), our framework explicitly defines them as a Gaussian vector that is independent of the limiting distributions $Z^{h}$ within the Netsor program. This independence facilitates the computation of the limiting distribution of the attention outputs, as it allows us to treat the attention scores separately from the other variables in the Netsor program.
@@ 3.2 Statement
As a preparation, we first introduce the results of the convergence of the Netsor program without an attention layer. The following theorem is a slight simplification of Theorem 5.4 in Yan19b . [^2] Yan19b considers controlled functions instead of pseudo-Lipschitz functions, the former of which is a generalization of pseudo-Lipschitz functions.
Fact 3.1 (Netsor Master Theorem Yan19b ).
@@
It may be noted that in the statement of Theorem 3.1, the boundedness of $\phi^{i,j}$ as in Assumption 3.1 is not essential, and we can drop this condition. However, we keep it for simplicity of the proof. Also, note that in this statement, $\mathring{p}_{1},\dots,\mathring{p}_{r}$ are random variables with positive variance, and the convergence in distribution ${\to}^{d}$ described in the statement refers to convergence to the distribution of them. Unlike the conventional master theorem derived from Tensor Programs Yan19a ; Yan19b , this theorem simultaneously characterizes the randomness of the ordinary variables like $g^{1},\dots,g^{m}$, along with the effects of finite-dimensional random variables $\mathring{p}_{1},\dots,\mathring{p}_{r}$.
@@
The following corollary is an analogue of Theorem A.5 in Yan20b .
@@
We present specific multi-head attention results as an example of the application of Corollary 3.2. This example implies that the output of the attention mechanisms is described by a non-Gaussian distribution. Specifically, when $y^{i}$ is an output of a multi-head attention layer as $\mathrm{MultiHead}_{i\cdot}$ defined in Eq. (4), $(Z^{y^{1}},\dots,Z^{y^{s}})$ follows a hierarchical Gaussian distribution whose conditioning $(\mathring{p}_{1},\dots,\mathring{p}_{r})=\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ itself is a random variable, causing $(Z^{y^{1}},\dots,Z^{y^{s}})$ to converge to a heavy-tailed non-Gaussian distribution overall. We provide the details as follows:
@@
It is well-known in probability theory that hierarchical mixtures of Gaussian components typically lead to distributions with heavier tails than standard Gaussians. This observation highlights a significant challenge for theoretical analysis, since many existing frameworks, such as NNGP and Tensor Programs, are developed under regimes where the limiting distributions are Gaussian (or can be expressed as Gaussian components with correction terms). Extending these tools to rigorously handle such heavy-tailed non-Gaussian behaviors remains an important open problem, and we believe our work provides a concrete starting point for such future developments.
@@
Our result in Example 3.1 has a close relationship with the result of HBSDN (20). First, the limiting distribution of the attention scores in our framework perfectly matches the Gaussian distribution in their infinite-head limit. Second, the variance of our finite-head non-Gaussian output distribution theoretically matches the variance of their infinite-head Gaussian output distribution. This alignment suggests that while the infinite-head limit correctly captures the second-order statistics of the output, our finite-head analysis provides a more complete picture by characterizing the non-Gaussian aspects of the distribution that are not captured by second-order statistics alone.

### v3

@@ 1 Introduction
A useful approach to understanding the complex probabilistic behavior of neural networks is through the study of parameter distributions in the infinite-width limit. Notable examples include the neural network Gaussian process (NNGP) lee2018deep; hron2020infinite, which approximates the limit of stochastic parameter distributions with Gaussian processes; the neural tangent kernel (NTK) jacot2018neural, which represents the model near the initial value with kernel functions; mean field theory, which describes the update of parameter distributions mei2018mean; and the Tensor Program, developed by yang2019scaling; yang2020tensor2; yang2021tensor2b; yang2020tensor3; yang2021tensor4; yang2023tensor4b; yang2022tensor5; yang2024tensor6, which is a general probabilistic analytical framework that unifies the representation of the infinite-width limit for a wide range of neural architectures and multiple layers. These methods provide probabilistic models that closely approximate the complex phenomena of neural networks.
One challenge in the studies is to properly represent the attention mechanisms used in Transformers vaswani2017attention, which frequently appear in recent large-scale architectures devlin2019bert; achiam2023gpt; zhao2023survey. Unlike ordinary multi-layer perceptrons, an attention layer has interactions between query and key variables by multiplication, which makes the infinite-width limit distribution considerably more complex as shown in hron2020infinite, for example. To avoid this difficulty, the existing studies have mainly limited themselves to two special settings to compute the parameter distribution of the limit: (i) Infinite-head regime: hron2020infinite considers the NNGP for the attention mechanism with infinite heads, resulting in a Gaussian approximation. (ii) $1/n$-scaling regime: the Tensor Programs yang2019tensor1 approximate the multiplication by changing the scale of the input variables for attention layers from $1/\sqrt{n}$ to $1/n$, where $n$ is a dimension of variables. However, these simplifications compromise the expressiveness and structure of actual attention mechanisms (and consequently, of Transformers). Specifically, the limit distribution at the infinite-head is not an effective approximation of the actual attention mechanisms because it differs significantly from the finite-head case. Also, the $1/n$-scaling regime makes all similarity scores converge to zero in the infinite-width limit, making the model equivalent to not measuring similarity between key and query vectors. Therefore, the limit parameter distribution of the attention mechanism is still under development.
@@ 1.1 Related Works
The concentration of measures in neural networks plays a fundamental role in both theoretical analysis and practical applications of machine learning. Initially, studies in this area aimed to characterize the feedforward signal propagation in wide neural networks with random weights, inspired by statistical mechanics Amari1977neural; sompolinsky1988chaos. The outputs of such random neural networks converge to a Gaussian process, and the computation of signal propagation reduces to the composition of kernel functions of Gaussian processes utilized in machine learning neal1996bayesian; williams1996computing; daniely2016toward; lee2018deep; matthews2018gaussian. They are often referred to as the NNGP. Since the NNGP kernel captures intrinsic inductive biases of architectures, its prediction performance correlates well with trained networks across different architectures lee2018deep. Moreover, one can interpret NNGP as a network with random initialization of optimization, which naturally leads to quantitative insight into desirable weight scales to avoid exploding/vanishing signal and gradient problems poole2016exponential; schoenholz2017deep.
@@
Depending on the network architecture, various types of kernels can be obtained, not only for fully-connected neural networks, but also convolutional neural networks (CNNs) novak2019bayesian, skip connections yang2017mean, naive or gated recurrent neural networks (RNNs) chen2018dynamical; yang2019tensor1, and more yang2019mean; gao2023wide. While some classical works assumed random Gaussian weights generated in an i.i.d. manner, recent research has shown that the same NNGP can be derived even in networks with non-Gaussian golikov2022non or weakly correlated weights saada2024beyond. The NNGP can also be obtained for a network with a narrow bottleneck layer sandwiched between wide layers agrawal2020wide.
@@
Two pioneering studies have investigated the NNGP of self-attention layers yang2019tensor1; hron2020infinite. Initially, yang2019tensor1 pointed out that an unconventional scaling factor of $1/n$ in the softmax function enables straightforward NNGP evaluation. Greg Yang has introduced the theoretical framework of Tensor Programs yang2019scaling; yang2020tensor2; yang2020tensor3, systematically composing kernel functions, including NNGPs, for modern neural networks. Transformers with the $1/n$-scaling fall within the scope of the applicability of Tensor Programs. However, the $1/n$-scaling is rarely used in theoretical or practical contexts, leaving the more realistic $1/\sqrt{n}$-scaling unresolved. To attack this problem, hron2020infinite analyzed self-attention by varying the number of heads. They numerically demonstrated non-Gaussian behavior emerging in the single-head case due to stochasticity and correlations within the attention matrix, even at infinite embedding dimensions. They further showed that if we take the infinite limit of the number of heads, Gaussian behavior emerges in the self-attention output and defines an NNGP kernel termed infinite attention. Although this infinite attention empirically improved performance on certain NNGP regression benchmarks, its suitability as a theoretical foundation for realistic self-attention remains uncertain. This is because practical attention implementations typically use only 1-128 heads everett2024scaling, far fewer than the embedding dimension.
@@
Beyond the classical NNGP and Tensor Programs analyses, several recent works have examined Transformers under the standard $1/\sqrt{n}$ scaling and related asymptotic or dynamical regimes. dinan2023effective analyzed Transformers by tracking the first two moments (the kernel) of the signal to characterize propagation at initialization and during training. cowsik2024geometric applied mean-field theory to characterize the edge of chaos via forward and backward signal propagation, assuming Gaussianity of the QK product. bordelon2024infinite employed dynamical mean field theory to study training dynamics under various infinite limits, including infinite width, heads, and depth, identifying parameterizations that ensure stable feature learning over time. noci2023shaped modeled signal evolution with a stochastic differential equation, which requires modifying the softmax function.
@@ 2.1 Notation and Setup of Neural Networks
We define notation for standard neural networks, excluding the attention mechanism. In particular, we adopt notations inspired by the framework of Tensor Programs yang2019scaling; yang2019tensor1, which allow us to describe a broad class of neural network architectures.
@@
This type of Tensor Program is called Netsor, and it covers a broad range of neural network architectures, including a perceptron layer, a convolutional layer, a recurrent layer, and many others yang2019scaling; yang2019tensor1.
@@
Here, the common scaling $1/\sqrt{n}$ in the definition of $p_{i,j}^{(a)}$ is a key issue in this study. [^1] Our analysis focuses on the $1/\sqrt{n}$-scaling, a choice motivated by its prevalence in both practical Transformer implementations and the theoretical literature (e.g., hron2020infinite; dinan2023effective; cowsik2024geometric). While some analyses have raised questions about its stability yang2022tensor5; bordelon2024infinite, the precise conditions for stable training remain an active area of research and may depend on factors often abstracted away in simplified limits, such as the number of tokens or data-specific statistics. For instance, recent work has shown that data statistics can significantly alter optimal scaling rules hayou2025optimal. Applying the row-wise softmax function to the score matrix, we define
@@
This setup of weight matrices is common in practice (see glorot2010understanding for summary) and also identical to that in the existing Tensor Programs yang2019scaling; yang2019tensor1; yang2020tensor3.
@@ 2.2 Distribution of Attention Mechanism in Previous Setup
We discuss the existing challenges on characterizing the output distribution of the attention layer, which exhibits non-Gaussian behavior. To derive this distribution, two regimes are typically considered: the choice of scaling in the similarity computation, and the number of heads.
@@
yang2019tensor1 studies an attention layer whose scaling term $1/\sqrt{n}$ in the scaled dot-product of Eq. (2) is replaced by $1/n$. In this regime, the Tensor Programs framework can show that the outputs of the attention layer converge to a Gaussian distribution: as $n\to\infty$, it holds that
@@
with some $\kappa>0$ (see Appendix A and Theorem E.8 in yang2019tensor1 for details). Intuitively, in the $1/n$-scaling regime, the dot-product score $p_{i,j}^{(a)}$ converges to 0 for every pair $(i,j)\in[s]^{2}$ and head $a\in[H]$, which simplifies the non-Gaussian behavior of the attention outputs.
@@
hron2020infinite considers the regime in which many dimensionalies diverge to infinite, then shows the convergence of the attention outputs to Gaussian. Specifically, it is shown that as $n,H\to\infty$, it holds that
@@
with some $\kappa^{\prime}>0$ (see Theorem 3 in hron2020infinite). The variance $(\kappa^{\prime})^{2}$ is described by a covariance of the nonlinearly transformed dot-product score $p_{i,j}^{(a)}$. In the result, by letting the number of heads $H$ grow to infinite, the complex non-Gaussian effects from the dot-product score are smoothed out, yielding a convenient Gaussian distribution.
While these approximations are analytically appealing, they have notable limitations. In practice, Transformers typically use $1/\sqrt{n}$-scaling (see Equation (1) in vaswani2017attention) and a finite number of heads, resulting in the frequently observed non-Gaussian behaviors of attention layers. Although the above regimes simplify the behavior into a Gaussian form for tractability, a more precise theory is needed to capture the true behavior of attention layers in practice.
@@ 3.1 Limiting Distribution
In Definition 3.1, the formulations for the limiting random variables $Z^{h}$ associated with initial vectors, MatMul outputs, and Nonlin outputs are developed by the existing Tensor Programs framework (yang2019tensor1). Specifically, the limiting distribution of initial vectors follows that given in Section 2.1; variables generated by MatMul converge to Gaussian distribution; and variables generated by Nonlin are nonlinear transformations of the corresponding limiting distributions.
The distinct aspect of our analysis lies in the treatment of the limiting distribution for the scalar dot-products $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, which represent pre-softmax attention scores. While the marginal Gaussian distribution of these scores is consistent with that in the infinite-head limit (see Theorem 3 in hron2020infinite), our framework explicitly defines them as a Gaussian vector that is independent of the limiting distributions $Z^{h}$ within the Netsor program. This independence facilitates the computation of the limiting distribution of the attention outputs, as it allows us to treat the attention scores separately from the other variables in the Netsor program.
@@ 3.2 Statement
As a preparation, we first introduce the results of the convergence of the Netsor program without an attention layer. The following theorem is a slight simplification of Theorem 5.4 in yang2019tensor1.222yang2019tensor1 considers controlled functions instead of pseudo-Lipschitz functions, the former of which is a generalization of pseudo-Lipschitz functions.
Fact 3.1 (Netsor Master Theorem yang2019tensor1).
@@
It may be noted that in the statement of Theorem 3.1, the boundedness of $\phi^{i,j}$ as in Assumption 3.1 is not essential, and we can drop this condition. However, we keep it for simplicity of the proof. Also, note that in this statement, $\mathring{p}_{1},\dots,\mathring{p}_{r}$ are random variables with positive variance, and the convergence in distribution ${\to}^{d}$ described in the statement refers to convergence to the distribution of them. Unlike the conventional master theorem derived from Tensor Programs yang2019scaling; yang2019tensor1, this theorem simultaneously characterizes the randomness of the ordinary variables like $g^{1},\dots,g^{m}$, along with the effects of finite-dimensional random variables $\mathring{p}_{1},\dots,\mathring{p}_{r}$.
@@
The following corollary is an analogue of Theorem A.5 in yang2020tensor3.
@@
We present specific multi-head attention results as an example of the application of Corollary 3.2. This example implies that the output of the attention mechanisms is described by a non-Gaussian distribution. Specifically, when $y^{i}$ is an output of a multi-head attention layer as $\mathrm{MultiHead}_{i\cdot}$ defined in Eq. (4), $(Z^{y^{1}},\dots,Z^{y^{s}})$ follows a hierarchical Gaussian distribution whose conditioning $(\mathring{p}_{1},\dots,\mathring{p}_{r})=\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ itself is a random variable, causing $(Z^{y^{1}},\dots,Z^{y^{s}})$ to follow a non-Gaussian distribution overall. We provide the details as follows:
@@
The non-Gaussianity of the limiting distribution of the attention layer highlights a significant challenge for theoretical analysis, since many existing frameworks, such as NNGP and Tensor Programs, are developed under regimes where the limiting distributions are Gaussian (or can be expressed as Gaussian components with correction terms). Extending these tools to rigorously handle such non-Gaussian behaviors remains an important open problem, and we believe our work provides a concrete starting point for such future developments.
@@
Our result in Example 3.1 has a close relationship with the result of hron2020infinite. First, the limiting distribution of the attention scores in our framework perfectly matches the Gaussian distribution in their infinite-head limit. Second, the variance of our finite-head non-Gaussian output distribution theoretically matches the variance of their infinite-head Gaussian output distribution. This alignment suggests that while the infinite-head limit correctly captures the second-order statistics of the output, our finite-head analysis provides a more complete picture by characterizing the non-Gaussian aspects of the distribution that are not captured by second-order statistics alone.
Remark 3.3.
We can show that the limiting distribution $Z^{y^{i}}$ is sub-Gaussian, provided that bounded nonlinearities are used for the attention scores (e.g., the softmax function as in Example 3.1). In other words, although $Z^{y^{i}}$ cannot be approximated by a single Gaussian distribution with mean $\mathbb{E}(Z^{y^{i}})$ and variance $\mathrm{Var}(Z^{y^{i}})$, its tail probability decays at a Gaussian rate, possibly with a larger variance parameter. A detailed discussion is provided in Appendix E.